OpenStudy (anonymous):

An automobile company has 10 choices of body styles and 8 choices of engines. How many combinations are available? A. 18 B. 40 C. 80 D. 120 I keep getting 45 >_>

OpenStudy (anonymous):

(8)(10) = 80

OpenStudy (anonymous):

Yeah that is right @swagmaster47

OpenStudy (anonymous):

I figured it out, Im a whole new kind of special haha!

OpenStudy (anonymous):

Need any other help @Ratedover

OpenStudy (anonymous):

What is the equation of a parabola with a vertex at (0, 0) and a directrix x = -4? XD

OpenStudy (anonymous):

Hmmmmmm?

OpenStudy (cj49):

x = -4 -p = -4 OR p = 4 and 4p = 16 y^2 = 16x OR x = (1/16)y^2

OpenStudy (anonymous):

:o THATS WHAT I GOT YEAH I GET IT! thanks :)

OpenStudy (cj49):

np...c:

OpenStudy (anonymous):

@cj49 I got this y=-(1/20)x^2 for this one, What is the equation of the parabola with a focus at (0, -5) and directrix y = 5?

OpenStudy (cj49):

correct

OpenStudy (anonymous):

You the bomb diggity! haha thanks

OpenStudy (cj49):

lol thnx xD